OpenStudy (anonymous):

y=sqrt(1+4sinx) (0,1) find the tangent to the curve at given point

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 2\cos(x) }{ \sqrt{1+4\sin(x)} }\]